The South Division of Wiig Company reported the following data for the current year. Sales Variable costs Controllable fixedNeed help with question B please
Answer #1

b) Calculate ROI

Return on investment
Action 1

(367000+99000)/5000000 = 9.32%

Action 2 (367000+150000)/5000000 = 10.34%
Action 3 367000/4850000 = 7.57%